OpenStudy (anonymous):

sin2x = 16/25 how do i get the value for x?

OpenStudy (anonymous):

is it \(\sin(2x)=\frac{16}{25}\) or \[\sin^2(x)=\frac{16}{25}\]?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

then the only thing i would know to do is write \[2x=\sin^{-1}(\frac{16}{25})\] and use a calculator to solve it